# Navigating in the Dark: A Multimodal Framework and Dataset for Nighttime Traffic Sign Recognition

### v1

@@ Abstract
Traffic signboards are vital for road safety and intelligent transportation systems, enabling navigation and autonomous driving. Yet, recognizing traffic signs at night remains challenging due to visual noise and scarcity of public nighttime datasets. Despite advances in vision architectures, existing methods struggle with robustness under low illumination and fail to leverage complementary mutlimodal cues effectively. To overcome these limitations, firstly, we introduce INTSD, a large-scale dataset comprising street-level night-time images of traffic signboards collected across diverse regions of India. The dataset spans 41 traffic signboard classes captured under varying lighting and weather conditions, providing a comprehensive benchmark for both detection and classification tasks. To benchmark INTSD for night-time sign recognition, we conduct extensive evaluations using state-of-the-art detection and classification models. Secondly, we propose LENS-Net, which integrates an adaptive image enhancement detector for joint illumination correction and sign localization, followed by a structured multimodal CLIP-GCNN classifier that leverages cross-modal attention and graph-based reasoning for robust and semantically consistent recognition. Our method surpasses existing frameworks, with ablation studies confirming the effectiveness of its key components. The dataset and code for LENS-Net is publicly available for research [^1] https://adityamishra-ml.github.io/INTSD/.
@@ 1 Introduction
Table 1: Overview of traffic sign datasets. TT100K has only 10000 images with signboard in it. §video-frames covering only 15,630 unique signs. 45 classes have more than 100 examples. The MVD data set only contains front and back signboard classes. signifies if the dataset contains more than 50% night-time images.
Dataset Images Classes Signs Attributes Region Boxes Unique Heavily Nightime
INTSD 6004 41 14,044 tampered, blur, glare, India $\checkmark$ $\times$ $\checkmark$
@@
Traffic signboards play a critical role in Advanced Driver Assistance Systems (ADAS), autonomous driving, navigation, and smart city systems. However, most existing traffic sign datasets [42, 54, 9] predominantly feature daytime imagery, leaving night-time recognition underexplored. To mitigate this gap, prior studies have employed generative translation or domain adaptation techniques such as CycleGAN [53] and CyCADA [18] to synthesize night-time scenes from daytime data. Yet, these approaches often introduce unrealistic artifacts and fail to capture complex photometric effects (such as glare, reflections, and sensor noise), resulting in a persistent domain gap that hinders robustness in real-world deployment [41, 2, 35]
To address this limitation, we present INTSD, a real-world Indian Nighttime Traffic Sign Dataset. The INTSD covers a wide range of geographical area across India and features 14k+ fully annotated instances, in diverse lighting and weather conditions. It also introduces four additional classes that are frequently misclassified by detection models but are absent in existing benchmarks. The inclusion of natural noise factors such as glare and motion blur, makes it a new standard for traffic sign detection and classification tasks.
@@
Recent advances in Graph Convolutional Neural Networks (GCNNs) have shown significant success in image classification [49, 40] and other domain classification tasks [5, 32, 30], largely due to their inherent ability to model spatial representations. Building on this, along with the proven ability of multimodal learning to fuse complementary data sources for more robust classification [36], we propose LENS-Net, Learned Enhancement & Nodal-Semantic Network for TSR. It is based on first detect and then classify principle and features: (i) an adaptive image enhancement based detector and (ii) a structured multimodal CLIP-GCNN based classifier. On INTSD, LENS-Net achieves 92.56% mAP@50 for signboard detection and macro-averaged precision of 78.89% for classification.
@@
    We present the first nighttime diverse traffic sign dataset with 6004 images across India. The dataset spans over 14K fully annotated images, covering 41 traffic signboard classes.
@@ 2 Literature Review
A wide range of traffic sign datasets support detection and recognition research (see Table 1). Early benchmarks such as GTSRB [44] and GTSDB [43] focused on well-lit, structured settings, while later datasets like SwedishTS [21], BelgiumTS [31], RTSD [42], and TT100K [54] expanded geographic and environmental diversity. Large-scale datasets including MVD [34], BDD100K [52], and MTSD [9] introduced complex contextual challenges. Yet, night-time datasets remain scarce, motivating INTSD, a dedicated benchmark for low-light traffic sign detection and recognition.
Image Enhancement: The night-time conditions adds noise and taint the computer vision tasks. Prior studies [48] have addressed this through image enhancement techniques aimed at improving visibility before detection. Recent works [29, 4] proposed adaptive and restoration-based frameworks that jointly optimize enhancement and detection., while zero-reference mapping [14] and illumination-aware methods [45] also show notable gains in nighttime perception. Motivated by these advances, our LENS-Net detector integrates adaptive image enhancement, enabling end-to-end illumination correction tuned for sign localization.
@@
Classification: Following detection, classification remains a fine-grained visual challenge, as many signs differ only in subtle symbols or numerals (e.g., speed limit 30 vs. 80). CNN-based classifiers have long dominated this space, outperforming earlier hand-crafted pipelines [54, 15], while recent approaches use attention [1, 51] and transformer-based [12] architectures to focus on discriminative regions. Such mechanisms are critical for distinguishing visually similar signs, especially under low-light or degraded conditions. In this work, our CLIP-GCNN based classifier builds upon this insight, combining cross-modal attention with graph-based reasoning for semantically coherent classification.
@@ 3 Dataset
The INTSD spans six districts across India, encompassing urban and rural settings and varying environmental conditions such as weather, illumination, motion blur, and glare. Each image contains multiple sign instances, ensuring dense spatial coverage. Figure 1 illustrates representative examples and their geographic distribution. The INTSD also includes challenging scenarios in which signboards are partially occluded by foliage, obscured by other signs, partially out-of-frame, or form part of larger signboards (examples in Appendix 7).
Figure 1: Geographical coverage and representative nighttime traffic sign examples from the INTSD.
@@ 3.1 Data Collection and Curation
The images are primarily extracted from videos recorded at 30fps using a low-budget ($\approx 60$ USD) action camera (AKASO Ek7000), supplemented with images from various smartphones (Samsung and IPhone), comprising a total of 6004 nighttime images. This mix results in images of varying quality, which we organize into high-quality (smartphone) and low-quality (action camera) subsets. We employed a semi-automated pipeline, iteratively training a YOLOv8 model, to efficiently segregate frames (sampling one frame per second) containing signs from those without signs.
Additionally, INTSD also includes 1,012 day images captured from the same routes. We limit the number of daytime images as they are well-represented in prior Indian datasets [50]. The structured organization (day/night and low/high-quality) makes INTSD useful for image enhancement and for day-to-night image translation tasks for traffic scenes.
@@
3.2 Annotation Process
Annotating traffic signboards is a time-consuming process as it involves selecting and approving images, marking traffic signs with bounding boxes, and assigning appropriate class label from the predefined Indian MoRTH taxonomy [^3] https://morth.nic.in/sites/default/files/road_safety_books.pdf. To increase efficiency, we divide this process into two sequential stages: (i) Manual annotation: We manually annotate images using the VGG Image Annotator (VIA) [8]. Following the protocol of MTSD [9], INTSD includes bounding boxes for all traffic-related sign types (e.g., directional, mandatory, informative) and assigns a corresponding class label to each. (ii) Quality checks: All annotations in INTSD underwent a strict, triple-verification protocol. Detection errors were minimal: 3.12% false negatives were identified and manually added. A 0.47% labeling error rate observed during the second verification was completely eliminated after the final review.
Additional Attributes. In addition to the standard traffic sign classes, we introduce four supplementary classes to address cases that are commonly confused by the TSR models and are not represented in existing datasets: (i) the tampered class includes signboards that are broken, rusted, vandalized, or covered with pamphlets; (ii) the unknown class comprises of signboards that cannot be mapped to any existing class due to occlusion, motion blur, glare, small size, or partial visibility within the frame; (iii) the advertisement class encompasses billboards and other large advertisement boards absent from current datasets; and (iv) the warning class contains signboards indicating diversions or other road warnings. INTSD also includes images affected by moderate motion blur and light glare, reflecting real-world conditions (visuals in Appendix 7).
3.3 Statistics
Class Distribution: The treemap in Figure 2 (left) illustrates the long-tail class distribution in INTSD. The top five classes: advertisement (4,204), direction (2,911), unknown (1,773), left_curve (855), and speed (851), cover most samples. While over half the classes have fewer than 100 instances. Rare but crucial signs such as airport, blow_horn, and cross_road are severely underrepresented. This imbalance challenges standard training and motivates the need for methods that learn robust representations across all classes.
Dataset Distributions: As illustrated in Figure 2 (right), the dataset statistics present several key challenges. Image sizes follow a bimodal distribution (golden), with distinct clusters around $\sqrt{area}\approx 1100$ and $\sqrt{area}\approx 3700$. Both the signboard sizes (red) and the number of signboards per image (green) exhibit a characteristic long-tail distribution. The majority of signboards are small, and most images contain only 1-4 signs, with an average of 2.32 instances per image. Geographical distribution of INTSD is provided earlier. We defer a detailed analysis of the modeling challenges posed by these distributions to Appendix 7.
@@ 4.1 LENS-Net Detector
YOLO models demonstrate strong capability in detecting small and densely packed objects, owing to advances in multi-scale feature aggregation and attention-based refinement mechanisms that preserve fine-grained spatial details. Motivated by these strengths, we adopt YOLOv8 as our signboard detector. However, the INTSD comprises images captured under highly varying illumination conditions. While some scenes exhibit good lighting due to street lamps and vehicle headlights, others are severely underexposed in the absence of such sources. This variability poses a significant challenge for a static data pre-processing pipeline. To address this issue, we wrap the detector’s backbone with a PreprocessWrapper that applies a learned per-image parameter prediction head and differentiable transforms, producing preprocessed images which are then fed for detection.
Figure 3: Overview of the proposed LENS-Net architecture. The LENS-Net detector (top) integrates adaptive image enhancement for joint illumination correction and sign localization. The LENS-Net classifier (bottom) employs a structured multimodal CLIP-GCNN module that leverages cross-modal attention and graph-based reasoning for sign classification.
Param Head: We introduce a small, convolutional parameter prediction head, $\Phi_{\text{head}}$, which takes a downsampled version of the normalized input image $\mathbf{I_{\text{norm}}}\in\mathbb{R}^{H\times W\times 3}$ and predict raw parameter vector $\mathbf{z}_{\text{raw}}\in\mathbb{R}^{3}$.
$\mathbf{z}_{\text{raw}}=\Phi_{\text{head}}(F_{\text{down}}(\mathbf{I_{\text{norm}}}))$ (1)
where $F_{\text{down}}$ is a bilinear resize operation to a fixed size. This vector $\mathbf{z}_{\text{raw}}$ is then scaled to produce three parameters: $\gamma$ (gamma-correction amount), $\alpha$ (brightness amount), and $\zeta$ (unsharp mask amount). The scaling is achieved via a tanh activation to map the output to (1,1)(-1,1), followed by a linear transformation to the parameter’s operational range ($p$):
$p=\frac{\tanh(z_{i})+1}{2}(p_{\text{max}}-p_{\text{min}})+p_{\text{min}},\>p\in\{\gamma,\alpha,\zeta\}$ (2)
Adaptive Image Enhancement: The $\mathbf{I_{\text{norm}}}$ is then transformed by this predicted set of parameters $p$. The differentiable transformation pipeline is given below:
  1. 1.
    Unsharp Masking: We first apply an unsharp mask to enhance edges. Mathematically, $\mathbf{I}_{U}=\mathbf{I_{\text{norm}}}+\zeta\cdot(\mathbf{I_{\text{norm}}}-G_{\sigma_{u}}(\mathbf{I_{\text{norm}}}))$, where $G_{\sigma_{u}}$ is a Gaussian blur operation with a fixed kernel sigma $\sigma_{u}$.
  2. 2.
    Gamma and Brightness Correction: The $\mathbf{I}_{U}$ is then corrected for contrast and brightness. $\mathbf{I}_{\text{final}}=\text{clamp}(\alpha\cdot\text{clamp}(\mathbf{I}_{U},\epsilon,1.0)^{\gamma},0.0,1.0)$, where $\epsilon$ is a small constant and clamp function is a standard mathematical operation, defined as: $\text{clamp}(x,a,b)=\max(a,\min(x,b))$. The inner clamp provides numerical stability for gradients, and outer clamp ensures a valid output data range.

### v2

# Learning Under Low Illumination: A Dataset and Algorithm for Traffic Sign Recognition

@@ Abstract
Traffic signboards are vital for road safety and intelligent transportation systems, enabling navigation and autonomous driving. Yet, recognizing traffic signs at night remains underexplored due to the scarcity of realistic public datasets capturing low-light degradations and distractor classes. Existing benchmarks are predominantly daytime and do not reflect challenges such as headlight glare, motion blur, sensor noise, and vandalized or ambiguous signage. To address these gaps, we introduce INTSD, a large-scale nighttime traffic sign dataset collected across diverse regions of India. INTSD contains street-level images spanning 41 traffic signboard classes, multiple distractor categories, and varied lighting and weather conditions. The dataset is designed to support both detection and fine-grained classification under realistic nighttime scenarios. To benchmark INTSD for nighttime sign recognition, we conduct extensive evaluations using state-of-the-art detection and classification models under standardized protocols. Additionally, we present LENS-Net, a strong baseline that integrates adaptive illumination-aware detection with multimodal semantic reasoning for robust nighttime sign classification. Experiments and ablations demonstrate the challenges posed by INTSD and establish competitive baselines for future research. The dataset and code for LENS-Net is publicly available for research [^1] https://adityamishra-ml.github.io/INTSD/.
@@ 1 Introduction
Table 1: Overview of traffic sign datasets. TT100K has only 10000 images with signboard in it. §video-frames covering only 15,630 unique signs. 45 classes have more than 100 examples. The MVD data set only contains front and back signboard classes. Unique means that the same physical signboard does not appear repeatedly in multiple images. Nighttime denotes datasets primarily collected under nighttime conditions.
Dataset Images Classes Signs Attributes Region Boxes Unique Nighttime
INTSD 11016 41 14,044 tampered, blur, glare, India $\checkmark$ $\times$ $\checkmark$
@@
CNTSSS [27] 4062 3 7687 $\times$ China $\checkmark$ $\times$ $\checkmark$
@@
Traffic signboards play a critical role in Advanced Driver Assistance Systems (ADAS), autonomous driving, navigation, and smart city systems. However, most existing traffic sign datasets [43, 58, 9] predominantly feature daytime imagery, leaving nighttime recognition underexplored. To mitigate this gap, prior studies have used generative translation or domain adaptation techniques such as CycleGAN [57] and CyCADA [17] to synthesize nighttime scenes from daytime data. Yet, these approaches often introduce unrealistic artifacts and fail to capture complex photometric effects (such as glare, reflections, and sensor noise), resulting in a persistent domain gap that hinders robustness in real-world deployment [42, 2, 36]. Additionally, prior study shows that generative translation to nighttime conditions do not accurately assess a model’s performance during the real nighttime scenario [27].
@@
To address this limitation, we present INTSD, a real-world Indian Nighttime Traffic Sign Dataset. The INTSD covers a wide range of geographical area across India and features 14k+ fully annotated instances, in diverse lighting and weather conditions. It also introduces four additional classes that are frequently misclassified by TSR models but are absent in existing benchmarks. The inclusion of natural noise factors such as glare and motion blur, makes it a new standard for traffic sign detection and classification tasks.
Recent advances in Graph Convolutional Neural Networks (GCNNs) have shown significant success in image classification [51, 41] and other domain classification tasks [5, 33, 30], largely due to their inherent ability to model spatial representations. Building on this, along with the proven ability of multimodal learning to fuse complementary data sources for more robust classification [37], we also propose LENS-Net, Learned Enhancement & Nodal-Semantic Network for sign recognition. It is based on first detect and then classify principle and features: (i) an adaptive image enhancement based detector and (ii) a structured multimodal CLIP-GCNN based classifier. On INTSD, LENS-Net achieves 92.56% mAP@50 for signboard detection and macro-averaged precision of 78.89% for classification.
@@
    We present the first diverse Indian nighttime traffic sign dataset with 11k+ images across India. The dataset spans over 14K fully annotated instances, covering 41 traffic signboard and supplementary classes.
@@
    We establish extensive baselines for detection and classification on INTSD, providing future research directions.
  3. 3.
@@ 2 Literature Review
A wide range of traffic sign datasets support detection and recognition research. Table 1 provides overview and comparison of publicly available benchmarks. Early benchmarks such as GTSRB [45] and GTSDB [44] focused on well-lit, structured settings, while later datasets like SwedishTS [21], BelgiumTS [31], RTSD [43], and TT100K [58] expanded geographic and environmental diversity. Large-scale datasets including MVD [35], BDD100K [56], and MTSD [9] introduced complex contextual challenges. Yet, nighttime datasets remain scarce, motivating INTSD, a dedicated benchmark for low-light traffic sign detection and recognition. CNTSSS [27] recently addressed nighttime detection with a limited set of traffic sign categories. INTSD extends this direction by enabling both detection and fine-grained classification by incorporating realistic distractor categories encountered in real-world driving.
Image Enhancement: The nighttime conditions adds noise and taint the computer vision tasks. Prior studies [50] have addressed this through image enhancement techniques aimed at improving visibility before detection. Recent works [29, 4] proposed adaptive and restoration-based frameworks that jointly optimize enhancement and detection, while zero-reference mapping [14] and illumination-aware methods [46] also show notable gains in nighttime perception. Motivated by these advances, our LENS-Net detector integrates adaptive image enhancement, enabling end-to-end illumination correction tuned for sign localization.
@@
Classification: Following detection, classification remains a fine-grained visual challenge, as many signs differ only in subtle symbols or numerals (e.g., speed limit 30 vs. 80). CNN-based classifiers have long dominated this space, outperforming earlier hand-crafted pipelines [58, 15], while recent approaches use attention [1, 55] and transformer-based [12] architectures to focus on discriminative regions. Such mechanisms are critical for distinguishing visually similar signs, especially under low-light or degraded conditions. In this work, our proposed CLIP-GCNN based classifier builds upon this insight, combining cross-modal attention with graph-based reasoning for semantically coherent classification.
@@ 3 Dataset
Traffic sign distributions vary widely across regions due to differences in urban, highway, and rural infrastructure [9]. To capture this diversity, the INTSD spans six districts across India, encompassing urban and rural settings and varying environmental conditions such as weather, illumination, motion blur, and glare. Each image contains multiple sign instances, ensuring dense spatial coverage. Figure 1 illustrates representative examples and their geographic distribution. The INTSD also includes challenging scenarios in which signboards are partially occluded by foliage, obscured by other signs, partially out-of-frame, or form part of larger signboards (examples in Section 1 of supplementary material).
Figure 1: Geographical coverage and representative nighttime traffic sign examples from INTSD. The dataset spans six districts and includes both rural and urban traffic signs captured under diverse lighting and weather conditions.
@@ 3.1 Data Collection and Curation
The images are primarily extracted from videos recorded at 30fps using a low-budget ($\approx 60$ USD) action camera (AKASO Ek7000), supplemented with images from various smartphones (Samsung, IPhone, and Google Pixel), comprising a total of 11,016 images. This mix results in images of varying quality, which we organize into high-quality (smartphone) and low-quality (action camera) subsets. Of these 11,016 images, 10,004 correspond to nighttime scenes, including 6,004 images containing traffic signboards and 4,000 nighttime images without signboards. We employed a YOLOv8 model to efficiently segregate frames (sampling one frame per second) containing signs from those without signs.
Furthermore, we organize the dataset based on (i) image quality and (ii) weather conditions. Images captured using action camera tend to be of lower quality, whereas those taken with smartphones are of higher quality. These subsets facilitate image-to-image translation tasks aimed at enhancing the visual quality of traffic or driving scenes. Besides nighttime images, we provide 1,012 day images captured from the same routes in both low-quality (from action camera) and high-quality (from smartphones) formats. These images can be utilized for day-to-night or night-to-day image translation tasks. We do not include a large set of daytime images due to their availability in prior Indian dataset [52].
3.2 Traffic Sign Class Taxonomy
We follow the official traffic sign taxonomy defined by the Ministry of Road Transport and Highways (MoRTH), Government of India [32], to map signboards to 41 class labels in INTSD. This set is selected in way that ensures there is no visual overlap within classes. It is important to highlight that the proposed taxonomy is inherently flexible, allowing for the incremental expansion of INTSD.
3.3 Annotation Process
Annotating traffic signboards is a time-consuming process as it involves selecting and approving images, marking traffic signs with bounding boxes, and assigning appropriate class labels to each bounding box. To increase efficiency, we divide this process into three sequential stages: (i) image selection, (ii) manual annotation, and (iii) quality checks.
Figure 2: Examples from the INTSD. Each image highlights a representative attribute or challenge present in the dataset (e.g., glare, blur, rain, tampered signs, etc.). The icon in the bottom-right corner indicates the image source (action camera or smartphone), while the label denotes the corresponding visual attribute.
Image Selection. We employ a YOLOv8 model to initially segregate frames extracted from recorded videos into two groups: images containing traffic signboards and those without signboards. Frames are sampled at a rate of one frame per second and subsequently verified through manual inspection to ensure quality and relevance. Since the images are derived from video sequences, we discard frames with severe glare or extreme motion blur where no traffic sign is discernible. However, INTSD intentionally retains images with light-to-moderate blur or glare when traffic signs remain visible, as such conditions frequently occur in real-world nighttime driving and introduce additional challenges for TSR.
Manual Annotation. We use the VGG Image Annotator (VIA) [8] to manually draw bounding boxes around traffic signboards and assign class labels to each corresponding box. All bounding boxes are subsequently verified for accuracy and completeness. INTSD includes bounding boxes for diverse traffic sign types, including directional, mandatory, informative, rural signs, highways signs, etc.
Distractor Classes. In addition to the standard traffic sign classes, we introduce four supplementary classes to address cases that are commonly confused by the TSR models and are not represented in existing datasets: (i) the tampered class includes signboards that are broken, rusted, vandalized, or covered with pamphlets; (ii) the unknown class comprises of signboards that cannot be mapped to any existing class due to occlusion, motion blur, glare, small size, or partial visibility within the frame; (iii) the advertisement class encompasses billboards and other large advertisement boards absent from current datasets; and (iv) the warning class contains signboards indicating diversions or other road warnings. INTSD also includes images affected by moderate motion blur and light glare, reflecting real-world conditions and enhancing the model’s robustness to such visual distortions. We provide the visual examples of the distractor classes and challenges present in the INTSD in Figure 2. Refer to the Section 1 of the supplementary material for more details and visuals.
Quality Checks. The annotations in INTSD are performed following a strict and standardized set of protocols to ensure high-quality labeling. Both detection and classification annotations are triple-verified to minimize the possibility of errors. Detection errors were minimal: 3.12% false negatives were identified and added. During the second verification of class labels, 0.47% labeling error rate was observed, which dropped to 0% after the third (final) verification stage.
3.4 Data Splits
We partition the dataset into five distinct splits. The five folds contain 1188, 1226, 1175, 1189, and 1226 images with 2133, 2134, 2129, 2133, and 2128 annotated signboard instances, respectively. Stratification preserves the class distribution in each fold, preventing rare classes from being underrepresented and maintaining a consistent distribution. The same holds true for the distractor classes. The exact fold definitions and annotation files are publicly released to ensure reproducibility and facilitate standardized benchmarking.
3.5 Statistics
The INTSD exhibits several characteristics that reflect the complexity of real-world nighttime sign recognition. We analyze key dataset statistics, including sensor diversity, class distribution, image resolution, signboard scale, object density, and geographical coverage. These statistics highlight the challenges posed by the dataset and motivate the need for robust detection and recognition methods.
Sensors diversity. INTSD is captured using a mixture of action cameras and smartphones, covering images from six different sensor types. This contrasts with datasets such as TT100K [58], which rely on images captured from a single sensor setup. The resulting sensor diversity introduces variations in image quality, color response, and noise characteristics, contributing to a more diverse and realistic evaluation benchmark.
Class Distribution. Figure 3 (left) provides the class distribution of the dataset. A primary characteristic of the dataset is a significant class imbalance, following a classic long-tail distribution. A small number of classes account for a large portion of the data. For instance, the top five most frequent classes: advertisement (4,204 instances), direction (2,911), unknown (1,773), left_curve (855), and speed (851), constitute a substantial majority of all labeled signs. Conversely, rare but crucial signs such as airport, blow_horn, and cross_road are severely underrepresented. This imbalance poses a significant challenge, as standard training methods can become biased towards the majority classes, leading to poor performance on the rare but equally important minority classes. This characteristic of the data strongly motivates the need for specialized methods that can learn robust representations for both the head and tail of the distribution.
@@ 3.5 Statistics
Image Sizes. Figure 3 (golden/right) depicts the bimodal distribution of the image sizes in the INTSD. It emphasizes that the images primarily come in two distinct size ranges: a large number of smaller images (around $\sqrt{area}\approx 1100$) and another group of much larger images (around $\sqrt{area}\approx 3700$).
Signboard Sizes. Figure 3 (red/right) illustrates the long-tail distribution of signboard sizes in INTSD. As evident, the majority of signboards are relatively small, with a progressively smaller proportion of larger signboards. This inherent size imbalance poses a significant challenge for detection models, as small signboards contain limited visual information and are more susceptible to being overlooked by standard detection architectures. Such a distribution highlights the importance of employing strategies like scale-aware feature extraction and data augmentation to mitigate the bias towards more prevalent size categories.
Signboards per Image. Figure 3 (green/right) also depicts a long-tail distribution of object density across the dataset with average of 2.32 signboards per image ($14044/6004$). Most images contain only a few signboards (1-4), whereas images with a higher number of signboards, corresponding to more cluttered scenes, are relatively less. This imbalance presents additional challenges for detection models, as densely populated scenes can introduce occlusions and complex spatial arrangements, making accurate localization and classification more difficult. Addressing this variation in object density requires models capable of robustly handling both sparse and cluttered scenarios.
Geographical Distribution. Figure 1 illustrates the geographical distribution of the images in our dataset. The images span six districts and encompass both rural and urban settings, as well as a diverse range of lighting conditions, including well-lit and poorly-lit environments. This diversity ensures that the dataset captures real-world variability.
Overall, these statistics demonstrate that INTSD captures diverse real-world conditions, making it a challenging benchmark for nighttime sign recognition.
Figure 4: Overview of the proposed LENS-Net architecture. The detector (top) integrates adaptive image enhancement for joint illumination correction and sign localization. The classifier (bottom) employs a structured multimodal CLIP-GCNN module that leverages cross attention and graph-based reasoning for sign classification.
@@ 4.1 LENS-Net Detector
YOLO models demonstrate strong capability in detecting small and densely packed objects, owing to advances in multi-scale feature aggregation and attention-based refinement mechanisms that preserve fine-grained spatial details. Motivated by these strengths, we adopt YOLOv8 as our signboard detector. However, the INTSD comprises images captured under highly varying illumination conditions. While some scenes exhibit good lighting due to street lamps and vehicle headlights, others are severely underexposed in the absence of such sources. This variability poses a significant challenge for a static data pre-processing pipeline. To address this issue, we wrap the detector’s backbone with a PreprocessWrapper that applies a lightweight parameter prediction head $\Phi_{\text{head}}$ that estimates per-image enhancement parameters (gamma, brightness, unsharp masking). These parameters are applied via differentiable transforms, producing preprocessed images which are then fed for detection. We refer the interested readers to the supplementary material (Section 3) for more details on $\Phi_{\text{head}}$ and adaptive image enhancement.
To ensure that preprocessing module learns meaningful transformations and avoids collapsing to trivial solutions, we employ two key strategies. (i) Stochastic Augmentation: During training, we apply a diverse set of strong, non-differentiable augmentations (photometric jitter, JPEG compression, Gaussian noise, blur, and small rotations) before the image is passed to the PreprocessWrapper. It forces $\Phi_{\text{head}}$ to learn to invert these degradations and normalize the image to a consistent domain for the detector. (ii) Preprocessing Loss: We apply an L2 regularization based loss ($\mathcal{L}_{\text{preproc}}$) that penalizes the predicted scaled parameters for deviating from their (no operation) default values: $\mathcal{L}_{\text{preproc}}\leftarrow\mathbb{E}\left[\sum_{i\in\{\gamma,\alpha,\zeta\}}\lambda_{i}(p_{i}-p_{\text{default},i})^{2}\right]$, where $\lambda_{i}$ ($\lambda_{i}\in(0,0.01)$) are per-parameter loss weights. This loss is added to the standard YOLOv8 detection loss, $\mathcal{L}_{\text{detect}}$, during training. Therefore, the net loss for detector is: $\mathcal{L}_{\text{total}}\leftarrow\mathcal{L}_{\text{detect}}+\mathcal{L}_{\text{preproc}}$.
LENS-Net detector is trained using a zero-one scheduler: (i) head-only optimization with frozen backbone, and (ii) joint end-to-end training. The detector’s bounding-box output ($\mathbf{I}_{\text{crop}}$) is forwarded to the classifier for recognition.
4.2 LENS-Net Classifier
After the detector localizes a traffic sign and outputs the corresponding cropped region, the classifier is tasked with accurately identifying the specific category. We design a tri-modal fusion classifier that combines: (i) robust, global semantic features from a frozen CLIP vision encoder, (ii) prompt-conditioned textual priors from CLIP text encoder, and (iii) structured attribute reasoning using a small graph. The complete architecture is detailed in Algorithm 1. The model processes the $\mathbf{I}_{\text{crop}}$ in two parallel branches that are fused for a final prediction.
Branch 1: Text-Image Fusion. We process $\mathbf{I}_{\text{crop}}$ to obtain the corresponding global image embedding $\mathbf{v}$, which is fused with semantic context derived from a set of pre-defined textual prompts $\mathbf{P}$. These prompts are encoded into embeddings $\mathbf{T}\in\mathbb{R}^{n(\mathbf{P})\times D}$ using the frozen CLIP text encoder (Algorithm 1, line 2-3). $\mathbf{v}$ and $\mathbf{T}$ are fused with cross-attention (Equation 1). This produces a prompt-conditioned image feature CA1, a refined global vector that already blends visual information with textual priors. CA1 is processed with residual connection (Res), Layer Normalization (LN), and position-wise feed-forward network (FFN) to produce the text-conditioned feature $\mathbf{f}_{\text{CA1}}$ (Equation 2).